# Universal logical operations with a dynamical qubit in Floquet code

### v1

@@ I Introduction
Figure 1: Implementation of the FBS code. a, Schematic of the quantum device for the FBS code, consisting of 9 data qubits (white circles labeled as $D_{i}$, where i = 1 to 9) and 12 ancilla qubits (orange circles for $Z$ ancillas and blue circles for $X$ ancillas). The four central qubits are alternately used as $X$ and $Z$ ancillas in different stabilizer rounds. Extra ancilla qubits (grey circles) are used for circuit compression and the construction of logical CNOT gate. Each pair of qubits is connected by a tunable coupler (H shape). b, Illustration of stabilizer measurements for the BS code (left) and the FBS code (right). The FBS code features period-4 stabilizer measurements with evolving logical operators $X_{\rm d}$ (purple lines) and $Z_{\rm d}$ (green lines) for the dynamical qubit in different rounds. In contrast, the logical operators $X_{\rm s}$ and $Z_{\rm s}$ are fixed and identical for both codes. Weight-2 Pauli measurements are used to check the parity of the corresponding data qubits in the $X$ (blue) and $Z$ (orange) bases. c, Cumulative distributions of errors for simultaneous single-qubit gates, CZ gates, ancilla qubit readout, and data qubit idle (with dynamical decoupling) during ancilla readout are depicted. The median error rates are 0.07%, 0.97%, 1.58%, and 1.43%, respectively, as indicated by the vertical lines.
@@
Quantum computers hold potential for solving computational problems that are provably intractable for classical systems, with transformative applications in quantum simulation [1], cryptography [2], and machine learning [3]. However, their practical realization is hindered by a fundamental challenge: quantum information is inherently fragile, susceptible to inevitable environmental noise and control inaccuracies. To address this challenge, quantum error correction (QEC) has emerged as a critical framework [4, 5, 6]. By redundantly encoding logical qubits into larger systems with many physical qubits, QEC detects and corrects errors through repeated syndrome measurements. Combined with fault-tolerant (FT) logical operations, QEC not only safeguards quantum information, but also lays the foundation for scalable and reliable quantum computation.
@@
A variety of QEC codes have been proposed, including surface code [7], color code [8], Bacon-Shor (BS) code [9], and quantum LDPC codes [10, 11]. Significant experimental progresses have been achieved in recent years across various platforms [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Notably, exponential error suppression has been demonstrated by scaling up a surface-code logical qubit using superconducting circuits [22]. While high-performance logical qubits can be realized by increasing the size of quantum processors, the code rate–the ratio of logical qubits to physical qubits–remains a crucial metric [11]. Low encoding efficiency can make FT quantum computing resource-intensive, especially for solid-state quantum devices limited to nearest-neighbor coupling.
@@
Recently proposed time-dynamical QEC, including Floquet codes [23, 24] and generalized time-dynamical code implementations [25, 26, 27, 28, 29], opens up new opportunities for fault-tolerant quantum computation. A recent work [30] has shown the flexibility in surface code realization using time-dynamical implementations. Unlike conventional stabilizer or subsystem codes that encode logical information in a fixed subspace, the Floquet codes allow logical information to evolve periodically over time. Additional dynamical logical qubits are encoded through an appropriate measurement schedule, thereby improving the code rate. During the periodic schedule of Pauli measurements, logical subspaces and operators adapt to stabilizer measurements while preserving the encoded information. Despite extensive theoretical study since its inception [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], the Floquet code has not yet been experimentally demonstrated.
@@
In this work, we implement the Floquet-Bacon-Shor (FBS) code [42] on a superconducting quantum processor, encoding one conventional static logical qubit and one dynamical logical qubit simultaneously within a $3\times 3$ array of data qubits. We demonstrate FT encoding and measurement of the two-qubit logical states, and stabilize these states using multi-round stabilizer measurements. We showcase FT Pauli gates and non-fault-tolerant (nFT) arbitrary rotation gates on the dynamical qubit. Furthermore, we achieve a logical Bell state fidelity of $75.9\%$ by entangling the dynamical and static logical qubits via a logical CNOT gate, with the gate fidelity assessed at $84.1\%$ through logical quantum process tomography (LQPT). With these demonstrated universal two-qubit logical operations, our results highlight the potential of Floquet codes for scalable and resource-efficient QEC.
@@ II Results
Figure 2: Encoding, stabilization, and measurement of two-qubit logical states in the FBS code. a, Simplified circuit for encoding, stabilizing and measuring two-qubit logical states. All physical qubits are initialized to the ground state before the circuit. Period-4 stabilizer measurements are alternately performed following the encoding circuit. b, Encoding circuits for logical Pauli states $|+,0\rangle$ and $|1,0\rangle$. BS stabilizer measurements are included in the latter case, with the dynamical qubit state post-selected to be $|0\rangle$. c, Example logical measurement circuits after stabilizer $S_{Z}^{\rm D}$: FT $ZZ$ measurement (left) and nFT $XZ$ measurement (right). Logical operators are indicated within the dashed-line boxes. Measurement outcomes of some data qubits are used to calculate the values of the $X$ (blue) and $Z$ (orange) stabilizers, enabling error detection. d, LQST fidelities of all 36 two-qubit Pauli states. The 16 states $\{|0\rangle,|1\rangle,|+\rangle,|-\rangle\}^{\otimes 2}$ are fault-tolerantly prepared and exhibit higher fidelity, as shown to the left of the dashed line. The blue portions indicate the fidelities without error detection, while the green portions highlight the fidelity improvement with error detection in the logical measurement. e, Experimentally measured (filled symbols) and simulated (open symbols) expectation values of $X_{\rm s}Z_{\rm d}$ operator versus stabilizer round, for the encoded states $|-,1\rangle$ (top) and $|-,0\rangle$ (bottom). The solid and dashed lines represent exponential fits. Each stabilizer round lasts 920 ns, comprising 720 ns for readout pulses and resonator photon depletion, and 200 ns for single-qubit and two-qubit gates. The cyan and green colors correspond to data without post-processing and data with measured error syndromes post-selected out, respectively. Due to the exponential decrease in the retained data during error detection, the analysis is limited to 16 rounds, with a total of 100,000 experimental shots. The insets show the error detection probabilities for the weight-6 stabilizers over 24 measurement rounds, with the final round calculated from the logical measurements.
@@
The BS code can be stabilized in a single measurement round using only weight-2 Pauli measurements, whose ordering is inconsequential [9], as shown in Fig. 1b. The four weight-6 stabilizers, $S_{X}^{\text{A}}$, $S_{Z}^{\text{B}}$, $S_{X}^{\text{C}}$, and $S_{Z}^{\text{D}}$, are derived by multiplying the outcomes of the corresponding Pauli measurements, enabling the detection and subsequent correction of any single-qubit error. Since these Pauli measurements do not always commute, the Hilbert space of the 9 physical qubits consists of 4 degrees of freedom fixed by the stabilizers, 1 degree for logical information, and 4 additional degrees corresponding to the gauge space, which does not encode logical information. By imposing a period-4 ordering on the weight-2 Pauli measurements, the gauge space dynamically encodes an additional logical qubit, thereby forming the FBS code. In each round, $4$ weight-$2$ Pauli measurements are performed, whose outcomes determine one weight-$6$ stabilizer and also influence the evaluation of dynamical operators (see Methods). The logical Pauli operators $X_{\text{d}}$ and $Z_{\text{d}}$ of the dynamically encoded qubit evolve periodically during the stabilizer measurements, as indicated by the purple and green lines in Fig. 1b. Both the static and dynamical logical qubits are stabilized over the 4 measurement rounds. Although the $3\times 3$ system size is insufficient for error correction of the dynamical qubit, the stabilizers extracted from the measurements are sufficient to detect any single-qubit error. Within the framework of error detection, an operation is considered FT if any single-qubit error produces a non-trivial syndrome and is thereby detectable [47].
@@
We perform period-4 stabilizer measurements to preserve the encoded two-qubit logical states. Specifically, the four stabilizers $S_{X}^{\rm A}$, $S_{Z}^{\rm B}$, $S_{X}^{\rm C}$, and $S_{Z}^{\rm D}$ are alternately performed to detect bit-flip or phase-flip errors in both the static and dynamical qubits. Each measurement result is compared to the result of the same stabilizer type from four rounds earlier, with the first four rounds compared to the encoding value. Errors during this process manifest as changes in the measurement outcomes, and the corresponding data are post-selected out in error detection. We present the preservation results for the encoded states $|-,0\rangle$ and $|-,1\rangle$ in Fig. 2e, showing the measured expectation values of the operator $X_{\rm s}Z_{\rm d}$. The logical error rates ($\varepsilon_{\rm exp}$) for these states, extracted by fitting the exponential decay of the expectation values, are reduced from 14.7% and 14.3% to 2.6% and 2.5%, respectively. This improvement corresponds to an increase in the logical lifetime $T_{\rm L}=\tau_{\rm exp}/2\varepsilon_{\rm exp}$ from 3.1 $\mu$s and 3.2 $\mu$s to 17.7 $\mu$s and 18.4 $\mu$[17], respectively. Here, $\tau_{\rm exp}=920$ ns is the duration of a stabilizer round. Numerical simulation results using the open-source package Stim [59], based on monitored gate and measurement errors, are also shown for comparison. The larger error rates observed in experiments may arise from leakage-induced correlated errors [60, 61, 62, 54, 63, 64, 58]. We measure an average leakage rate of 0.18% per round on the physical qubits using three-state readout at the end of the circuit (see Supplementary Information IV). In both the experimental and simulation data, round-dependent oscillations are observed. Specifically, higher expectation values occur every four rounds, corresponding to the completion of a full error detection cycle. Beyond error detection, error correction can be applied to the static qubit alone (see Supplementary Information VI). However, the effectiveness of error correction diminishes for the two-qubit states due to the small code distance of the dynamical qubit and the correlation between errors in the two logical qubits.
@@
The construction of arbitrary rotation gates is facilitated by the use of ancilla qubits. For logical rotations around the $Z_{\rm d}$-axis ($R_{Z}^{\varphi}$), a physical gate $R_{z}^{\varphi}$ is inserted on an ancilla qubit and placed between four physical CNOT gates acting on $D_{7}$ and $D_{8}$, as illustrated at the top of Fig. 3b. The entire circuit is then inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. Similarly, logical rotations around the $X_{\rm d}$-axis ($R_{X}^{\theta}$) are implemented using an $R_{z}^{\theta}$ gate on an ancilla qubit. The corresponding circuit is inserted between stabilizers $S_{X}^{\rm C}$ and $S_{Z}^{\rm D}$. We sweep the rotation angles $\varphi$ and $\theta$ and measure the dynamical qubit in three bases: $X_{\rm d}$, $Y_{\rm d}$, and $Z_{\rm d}$. The measurement results are shown in Fig. 3b, with filled (open) symbols representing data with (without) error detection. With error detection enabled, the lower amplitude in the $Y_{\rm d}$-direction is primarily due to the nFT nature of the logical $Y_{\rm d}$ measurement. Without error detection, the amplitude in the $Y_{\rm d}$-direction is lower than that in the $Z_{\rm d}$-direction for the $R_{X}^{\theta}$ rotation due to more phase-flip errors propagating through the nFT $R_{X}^{\theta}$ gate.
@@
Figure 4: Logical CNOT gate between the dynamical and static qubits. a, The CNOT circuit (top) is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$ in the overall circuit (bottom). For Bell state generation, the logical qubits are initialized to $|+,0\rangle$ in the encoding process. For LQPT, 16 linearly independent logical cardinal states: $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are encoded. b, Error-detected logical density matrix of the generated logical Bell state, obtained with four stabilizer rounds. The logical state fidelity is 75.9% with error detection and 38.8% without error detection. c, Extracted LPTM $R_{\rm exp}$ of the CNOT gate, with a process fidelity $F_{\rm p}=80.2\%$.
@@
Finally, we show that the dynamical and static qubits can be entangled using a logical CNOT gate. The specific CNOT circuit depends on the preceding stabilizer type. In our implementation, the CNOT circuit is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. This circuit involves 14 entangling gates acting on a subgroup of 3 data qubits and 2 ancilla qubits, as illustrated in Fig. 4a. The circuit is derived from using the static qubit operator $Z_{\rm s}=Z_{1}Z_{4}Z_{7}$ (which is equivalent to $Z_{\rm s}$ in Eq. 1 up to stabilizer $S_{Z}^{\rm D}$) to control the dynamical qubit operator $X_{\rm d}=X_{1}X_{7}$ after round $S_{Z}^{\rm B}$ (see Supplementary Information Section I for theoretical derivation). We note that the logical CNOT gate we implement is nFT because it allows errors to propagate between the 3 involved data qubits. The circuit is intentionally designed to involve only 3 data qubits to minimize error propagation.
@@
Applying the logical CNOT gate to the encoded state $|+,0\rangle$ yields a logical Bell state $\frac{1}{\sqrt{2}}(|0,0\rangle+|1,1\rangle)$. The error-detected logical density matrix $\rho_{\rm B}$ is depicted in Fig. 4b, with a fidelity of 75.9%. Without error detection, the corresponding fidelity is only 38.8%. We benchmark the logical CNOT gate fidelity using LQPT. 16 linearly independent two-qubit logical states $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are prepared in the encoding process. For each initial state, two-qubit LQST is performed at the end of the circuit. Due to the susceptibility of LQPT to state preparation and measurement (SPAM) errors, we characterize the input states using the same circuit without inserting the logical CNOT gate [65]. The logical Pauli transfer matrix (LPTM) $R_{\rm exp}$, which linearly matches the input states to output states, is calculated using the 16 input-output pairs (see Methods). The resulting LPTM with error detection is shown in Fig. 4b. The process fidelity is calculated by comparing $R_{\rm exp}$ with the ideal transfer matrix $R_{\rm ideal}$, yielding $\mathcal{F}_{p}={\rm Tr}(R_{\rm exp}\cdot R_{\rm ideal})=80.2\%$. This corresponds to a gate fidelity $\mathcal{F}_{g}=(\mathcal{F}_{p}/E+1)/(E+1)$ of 84.1%, where $E=4$ is the Hilbert space dimension. The fidelity calculation based on LQPT suffers from the low fidelity of encoded $|{-}i\rangle$ state on either qubit. We can also estimate the logical CNOT gate fidelity by dividing the average output state fidelity by the average input state fidelity without $|{-}i\rangle$, indicating a gate fidelity of 93.6%.

### v2

@@ I Introduction
Figure 1: Implementation of the FBS code. a, Schematic of the quantum device for the FBS code, consisting of 9 data qubits (white circles labeled as $D_{i}$, where i = 1 to 9) and 12 ancilla qubits (orange circles for $Z$ ancillas and blue circles for $X$ ancillas). The four central qubits are alternately used as $X$ and $Z$ ancillas in different stabilizer rounds. Extra ancilla qubits (grey circles) are used for circuit compression and the construction of logical CNOT gate. Each pair of qubits is connected by a tunable coupler (H shape). b, Illustration of stabilizer measurements for the BS code (left) and the FBS code (right). The FBS code features period-4 stabilizer measurements, during which the logical operators $X_{\rm d}$ (purple lines) and $Z_{\rm d}$ (green lines) for the dynamical qubit evolve across different rounds. In contrast, the logical operators $X_{\rm s}$ and $Z_{\rm s}$ for the static qubit remain fixed and are identical in both codes. Weight-2 Pauli measurements are used to check the parity of the corresponding data qubits in the $X$ (blue) and $Z$ (orange) bases. c, Cumulative distributions of simultaneous errors for 21 single-qubit gates, 32 CZ gates, ancilla qubit readout, and data qubit idle (with dynamical decoupling) during ancilla readout are depicted. The median error rates are 0.07%, 0.97%, 1.58%, and 1.43%, respectively, as indicated by the vertical lines.
@@
Quantum computers hold potential for solving computational problems that are provably intractable for classical systems, with transformative applications in quantum simulation [1], cryptography [2], and machine learning [3]. However, their practical realization is hindered by a fundamental challenge: quantum information is inherently fragile, susceptible to inevitable environmental noise and control inaccuracies. To address this challenge, quantum error correction (QEC) has been established as a critical framework [4, 5, 6]. By redundantly encoding logical qubits into larger systems with many physical qubits, QEC detects and corrects errors through repeated syndrome measurements. Combined with fault-tolerant (FT) logical operations, QEC not only safeguards quantum information, but also lays the foundation for scalable and reliable quantum computation.
@@
A variety of QEC codes have been proposed, including surface code [7], color code [8], Bacon-Shor (BS) code [9], and quantum LDPC codes [10, 11]. Significant experimental progresses have been achieved in recent years across various platforms [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Notably, exponential error suppression has been demonstrated by scaling up a surface-code logical qubit using superconducting circuits [23]. While high-performance logical qubits can be realized by increasing the size of quantum processors, the code rate–the ratio of logical qubits to physical qubits–remains a crucial metric [11]. Low encoding efficiency can make FT quantum computing resource-intensive, especially for solid-state quantum devices limited to nearest-neighbor coupling.
@@
Recently proposed time-dynamical QEC, including Floquet codes [24, 25] and generalized time-dynamical code implementations [26, 27, 28, 29, 30], opens up new opportunities for fault-tolerant quantum computation. A recent work [31] has shown the flexibility in surface code realization using time-dynamical approaches. Unlike conventional stabilizer or subsystem codes that encode logical information in a fixed subspace, the Floquet codes [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52] allow logical information to evolve periodically over time. Additional dynamical logical qubits are encoded through an appropriate measurement schedule, thereby improving the code rate. During the periodic schedule of Pauli measurements, logical subspaces and operators adapt to stabilizer measurements while preserving the encoded information.
@@
In this work, we implement the Floquet-Bacon-Shor (FBS) code [44] on a superconducting quantum processor, encoding one dynamical logical qubit and one conventional static logical qubit simultaneously within a $3\times 3$ array of data qubits. We demonstrate FT encoding and measurement of the two-qubit logical states, and stabilize these states using multi-round stabilizer measurements. We showcase FT Pauli gates and non-fault-tolerant (nFT) arbitrary rotation gates on the dynamical qubit. Furthermore, we achieve a logical Bell state with a fidelity of $75.9\%$ by entangling the dynamical and static logical qubits via a logical CNOT gate, with the gate fidelity assessed at $84.1\%$ through logical quantum process tomography (LQPT). With these demonstrated universal two-qubit logical operations, our results highlight the potential of Floquet codes for scalable and resource-efficient QEC.
@@ II Results
Figure 2: Encoding, stabilization, and measurement of two-qubit logical states in the FBS code. a, Simplified circuit for encoding, stabilizing and measuring two-qubit logical states. All physical qubits are initialized to the ground state before the circuit. Period-4 stabilizer measurements are alternately performed following the encoding circuit. b, Encoding circuits for logical Pauli states $|+,0\rangle$ and $|1,0\rangle$. BS stabilizer measurements are included in the latter case, with the dynamical qubit state post-selected to be $|0\rangle$. c, Example logical measurement circuits after stabilizer $S_{Z}^{\rm D}$: FT $ZZ$ measurement (left) and nFT $XZ$ measurement (right). Logical operators are indicated within the dashed-line boxes. Measurement outcomes of some data qubits are used to calculate the values of the $X$ (blue) and $Z$ (orange) stabilizers, enabling error detection. The notation $D_{147}$ denotes the three data qubits $D_{1}$, $D_{4}$, and $D_{7}$, and so on. d, LQST fidelities of all 36 two-qubit Pauli states. The 16 states $\{|0\rangle,|1\rangle,|+\rangle,|-\rangle\}^{\otimes 2}$ are fault-tolerantly prepared and exhibit higher fidelity, as shown on the left side (white background). The blue portions indicate the fidelities without error detection, while the green portions highlight the fidelity improvement with error detection in the logical measurement. e, Experimentally measured (filled symbols) and simulated (open symbols) expectation values of $X_{\rm s}Z_{\rm d}$ operator versus stabilizer round, for the encoded states $|-,1\rangle$ (top) and $|-,0\rangle$ (bottom). The solid and dashed lines represent exponential fits. Each stabilizer round lasts 920 ns, comprising 720 ns for readout pulses and resonator photon depletion, and 200 ns for single-qubit and two-qubit gates. The cyan and green colors correspond to data without post-processing (raw) and data with measured error syndromes post-selected out (detection), respectively. Due to the exponential decrease in the retained data during error detection, the analysis is limited to 16 rounds, with a total of 100,000 experimental shots. The insets show the error detection probabilities for the weight-6 stabilizers over 24 measurement rounds, with the final round calculated from the logical measurements.
@@
The BS code can be stabilized in a single measurement round using only weight-2 Pauli measurements, whose ordering is inconsequential [9], as shown in Fig. 1b. The four weight-6 stabilizers, $S_{X}^{\text{A}}$, $S_{Z}^{\text{B}}$, $S_{X}^{\text{C}}$, and $S_{Z}^{\text{D}}$, are derived by multiplying the outcomes of the corresponding Pauli measurements, enabling the detection and subsequent correction of any single-qubit error. Since these Pauli measurements do not always commute, the Hilbert space of the 9 physical qubits consists of 4 degrees of freedom fixed by the stabilizers, 1 degree for logical information, and 4 additional degrees corresponding to the gauge space, which does not encode logical information. By imposing a period-4 ordering on the weight-2 Pauli measurements, the gauge space dynamically encodes an additional logical qubit, thereby forming the FBS code. In each round, $4$ weight-$2$ Pauli measurements are performed, whose outcomes determine one weight-$6$ stabilizer and also influence the evaluation of dynamical operators (see Methods). The logical Pauli operators $X_{\text{d}}$ and $Z_{\text{d}}$ of the dynamically encoded qubit evolve periodically during the stabilizer measurements, as indicated by the purple and green lines in Fig. 1b. Throughout this paper, the initial operator definitions are consistent with stabilizer round A. Consequently, the first stabilizer round A does not change the encoded operators. Both the static and dynamical logical qubits are stabilized over the 4 measurement rounds. Although the $3\times 3$ system size is insufficient for error correction of the dynamical qubit, the stabilizers extracted from the measurements are sufficient to detect any single-qubit error. Within the framework of error detection, an operation is considered FT if any single-qubit error produces a non-trivial syndrome and is thereby detectable [53].
@@
We perform period-4 stabilizer measurements to preserve the encoded two-qubit logical states. Specifically, the four stabilizers $S_{X}^{\rm A}$, $S_{Z}^{\rm B}$, $S_{X}^{\rm C}$, and $S_{Z}^{\rm D}$ are alternately performed to detect bit-flip or phase-flip errors in both the static and dynamical qubits. Each measurement result is compared to the result of the same stabilizer type from four rounds earlier, with the first four rounds compared to the encoding value. Errors during this process manifest as changes in the measurement outcomes, and the corresponding data are post-selected out in error detection. We present the preservation results for the encoded states $|-,0\rangle$ and $|-,1\rangle$ in Fig. 2e, showing the measured expectation values of the operator $X_{\rm s}Z_{\rm d}$. The logical error rates ($\varepsilon_{\rm exp}$) for these states, extracted by fitting the exponential decay of the expectation values, are reduced from 14.7% and 14.3% to 2.6% and 2.5%, respectively. This improvement corresponds to an increase in the logical lifetime $T_{\rm L}=\tau_{\rm exp}/2\varepsilon_{\rm exp}$ from 3.1 $\mu$s and 3.2 $\mu$s to 17.7 $\mu$s and 18.4 $\mu$[17], respectively. Here, $\tau_{\rm exp}=920$ ns is the duration of a stabilizer round. Numerical simulation results using the open-source package Stim [65], based on monitored gate and readout errors, are also shown for comparison. The larger error rates observed in experiments may arise from leakage-induced correlated errors [66, 67, 68, 60, 69, 70, 64]. We measure an average leakage rate of 0.18% per round on the physical qubits using three-state readout at the end of the circuit (see Supplementary Information IV). In both the experimental and simulation data, round-dependent oscillations are observed. Specifically, higher expectation values occur every four rounds, corresponding to the completion of a full error detection cycle. Beyond error detection, error correction can be applied to the static qubit alone (see Supplementary Information VI). However, the effectiveness of error correction diminishes for the two-qubit states due to the small code distance of the dynamical qubit and the correlation between errors in the two logical qubits.
@@
The construction of arbitrary rotation gates is facilitated by the use of ancilla qubits (see Supplementary Information Section I for theoretical details). For logical rotations around the $Z_{\rm d}$-axis ($R_{Z}^{\varphi}$), a physical gate $R_{z}^{\varphi}$ is inserted on an ancilla qubit and placed between four physical CNOT gates acting on $D_{7}$ and $D_{8}$, as illustrated at the top of Fig. 3b. The entire circuit is then inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. Similarly, logical rotations around the $X_{\rm d}$-axis ($R_{X}^{\theta}$) are implemented using an $R_{z}^{\theta}$ gate on an ancilla qubit, and the corresponding circuit is inserted between stabilizers $S_{X}^{\rm C}$ and $S_{Z}^{\rm D}$. We sweep the rotation angles $\varphi$ and $\theta$ and measure the dynamical qubit in three bases: $X_{\rm d}$, $Y_{\rm d}$, and $Z_{\rm d}$. The measurement results are shown in Fig. 3b, with filled (open) symbols representing data with (without) error detection. With error detection enabled, the lower amplitude in the $Y_{\rm d}$-direction is primarily due to the nFT nature of the logical $Y_{\rm d}$ measurement. Without error detection, the amplitude in the $Y_{\rm d}$-direction is lower than that in the $Z_{\rm d}$-direction for the $R_{X}^{\theta}$ rotation, due to more phase-flip errors propagating through the nFT $R_{X}^{\theta}$ gate.
@@
Figure 4: Logical CNOT gate between the dynamical and static qubits. a, The CNOT circuit (top) is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$ in the overall circuit (bottom). For Bell state generation, the logical qubits are initialized to $|+,0\rangle$ in the encoding process. For LQPT, 16 linearly independent logical cardinal states: $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are encoded. b, Error-detected logical density matrix of the generated logical Bell state, obtained with four stabilizer rounds. The logical state fidelity is 75.9% with error detection and 38.8% without error detection. The wireframes indicate the ideal values. c, Extracted LPTM $R_{\rm exp}$ of the CNOT gate, with a process fidelity $F_{\rm p}=80.2\%$.
@@
Finally, we show that the dynamical and static qubits can be entangled using a logical CNOT gate. The specific CNOT circuit depends on the preceding stabilizer type. In our implementation, the CNOT circuit is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. This circuit involves 14 entangling gates acting on a subgroup of 3 data qubits and 2 ancilla qubits, as illustrated in Fig. 4a. The circuit is derived from using the static qubit operator $Z_{\rm s}=Z_{1}Z_{4}Z_{7}$ (which is equivalent to $Z_{\rm s}$ in Eq. 1 up to stabilizer $S_{Z}^{\rm D}$) to control the dynamical qubit operator $X_{\rm d}=X_{1}X_{7}$ after stabilizer $S_{Z}^{\rm B}$ (see Supplementary Information Section I for theoretical derivation). We note that the logical CNOT gate we implement is nFT because it allows errors to propagate between the 3 involved data qubits. The circuit is intentionally designed to involve only 3 data qubits to minimize error propagation.
@@
Applying the logical CNOT gate to the encoded state $|+,0\rangle$ yields a logical Bell state $\frac{1}{\sqrt{2}}(|0,0\rangle+|1,1\rangle)$. The error-detected logical density matrix $\rho_{\rm B}$ is depicted in Fig. 4b, with a fidelity of 75.9%. Without error detection, the corresponding fidelity is only 38.8%. We benchmark the logical CNOT gate fidelity using LQPT. 16 linearly independent two-qubit logical states $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are prepared in the encoding process. For each initial state, two-qubit LQST is performed at the end of the circuit. Due to the susceptibility of LQPT to state preparation and measurement (SPAM) errors, we characterize the input states using the same circuit without inserting the logical CNOT gate [71]. The logical Pauli transfer matrix (LPTM) $R_{\rm exp}$, which linearly matches the input states to output states, is calculated using the 16 input-output pairs (see Methods). The resulting LPTM with error detection is shown in Fig. 4c. The process fidelity is calculated by comparing $R_{\rm exp}$ with the ideal transfer matrix $R_{\rm ideal}$, yielding $\mathcal{F}_{p}={\rm Tr}(R_{\rm exp}\cdot R_{\rm ideal})=80.2\%$. This corresponds to a gate fidelity $\mathcal{F}_{g}=(\mathcal{F}_{p}/E+1)/(E+1)$ of 84.1% [72], where $E=4$ is the Hilbert space dimension. The fidelity calculation based on LQPT suffers from the low fidelity of encoded $|{-}i\rangle$ state on either qubit. We can also estimate the logical CNOT gate fidelity by dividing the average output state fidelity by the average input state fidelity without $|{-}i\rangle$, indicating a gate fidelity of 93.6%.